OpenStudy (anonymous):

On Saturday, a local hamburger shop sold a combined total of 357 hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Saturday?

OpenStudy (anonymous):

I developed an equation; 2x + x = 357. x = 119 so once you fill that into the equation it becomes; 2(119) + 119 = 357 238 + 119 = 357 x represents the hamburgers, so 119 hamburgers were sold on Saturday. :)

OpenStudy (anonymous):

thanks